OpenStudy (anonymous):

Add or Subtract SHOW WORK!!!!!!! 2 7/8 - 2/3

OpenStudy (anonymous):

i hate mixed numbers

hero (hero):

lol

hero (hero):

(2-0) + (7/8 - 2/3)

OpenStudy (lgbasallote):

nightmares

hero (hero):

(2-0) + (21/24 - 16/24)

OpenStudy (anonymous):

your first job is to write \(2\tfrac{7}{8}=\frac{7\times 2+8}{8}=\frac{16+7}{8}=\frac{23}{8}\)

OpenStudy (anonymous):

damn typo \[2\tfrac{7}{8}=\frac{8\times 2+8}{8}=\frac{16+7}{8}=\frac{23}{8}\]

hero (hero):

You don't need to change it to improper fraction

OpenStudy (anonymous):

your next job is to subtract \[\frac{23}{8}-\frac{2}{3}\] no but i do because i hate mixed numbers. i hate borrowing and carrying and all that junk

hero (hero):

Wow, you're a rare breed. I thought you liked doing things the easy way

hero (hero):

(2-0) + (21/24 - 16/24) 2 + 5/24

OpenStudy (anonymous):

but i suppose you are right, you can subtract \(\frac{7}{8}-\frac{2}{3}=\frac{7\times 3-8\times 2}{8\times 3}=\frac{21-16}{24}=\frac{5}{24}\)

OpenStudy (anonymous):

THANK YOU ALL FOR HELPING ME NOW I CAN GO TO BED!!!!XOXO ~Luv Savannah Nanami/ Little Japan school girl, na im from AMERIACA!!!!FREEDOM!!!